Read Activity 2-1 and Activity 2-2 in your lab man
Answer #1


while charging a capacitor the potential increase is given as

v = e*(1-e^-t/T)

T = time constant

1)


2.36 = 6*(1-e^-(4.16*10^-3/T1))


T1 = 8.32 ms

2)

3.79 = 6*(1-e^-(8.32*10^-3/T2))


T2 = 8.33 ms


3)

4.66 = 6*(1-e^-(12.48*10^-3/T3))


T3 = 8.32 ms

4)

5.19 = 6*(1-e^-(16.64*10^-3/T4))


T4 = 8.3 ms


average time constant = T = 8.32 ms   <<------answer

(b)

T = RC

C = T/R = (8.32*10^-3)/104 = 80 uF <<------answer